OpenStudy (compassionate):

You guys will argue over anything. 6÷2(1+2)

hartnn (hartnn):

already argued over this...and consensus was its 6

hartnn (hartnn):

sorry 9

OpenStudy (compassionate):

We are conversing, aren't we?

OpenStudy (kainui):

It's 9. Next question.

OpenStudy (compassionate):

I disagree that it's 9.

OpenStudy (abb0t):

Why is it 9 and not 13?

OpenStudy (compassionate):

Abb0t, demonstrate your calculations.

OpenStudy (karatechopper):

P E MD AS 6/2(1+2) 6/2(3) Left to right! MD 6/2=3 3(3)=9 My answer is 9

OpenStudy (abb0t):

\[\int\limits_{0}^{L/2}dx_1\int\limits_{0}^{L/2} dx_2|\Psi(x_1, x_2)|^2=\int\limits_{0}^{L/2}dx_1\int\limits_{0}^{L/2}dx_2|\Psi_1^2(x_1)\Psi_2^2(x_2)-\Psi_1(x_1)\Psi_1(x_2)\Psi_2(x_1)|\] Which is \[I_1I_2-I_{12}^2\] Where: \[I_1= \int\limits \Psi_1^2(x)dx= \frac{ 1 }{ 2 } \] and \[I_2= \int\limits \Psi_2^2(x)dx=\frac{ 1 }{ 2 } \] and \[I_{1,2} = \int\limits\limits \Psi_1(x) \Psi_2(x)dx = \pm \frac{ \sqrt{51} }{ 2 }i\] Resulting in 13

OpenStudy (karatechopper):

Demonstrate your calculation @Compassionate

OpenStudy (karatechopper):

WOAHH....@abb0t what did you bring in?!! I dont see any squiggles in the main problem!!

OpenStudy (abb0t):

pemdas is correct. It should be 9 though :) happy holidays all!

OpenStudy (karatechopper):

I said it was 9!! Happy that I am right!

OpenStudy (anonymous):

Ouch I have a headache... What in the world did you type @abb0t It's like in swahili...

OpenStudy (karatechopper):

Hahahah I couldn't even understand it.

OpenStudy (accessdenied):

Looks like stuff from Quantum Mechanics, at least that's the only time I ever saw Psi used like that... Although I don't know much about the subject. :p

OpenStudy (abb0t):

Particle in a box :)

OpenStudy (karatechopper):

eh?

OpenStudy (athenawolf):

it could also be 1, if you distributed.... :)

OpenStudy (abb0t):

Haha. My initial thought was 1, but I 4got about PEMDAS.

OpenStudy (unklerhaukus):

\[6÷2(1+2)=3(3)=9\]

OpenStudy (anonymous):

@Hero so you are saying you hate math

OpenStudy (anonymous):

9

OpenStudy (shubhamsrg):

can we really directly assume that its not 2*3 together there? with that, it comes to be 1 :/

OpenStudy (shubhamsrg):

i see..didnt know that..hmm..

OpenStudy (anonymous):

http://media.tumblr.com/tumblr_m2806ky0Wz1qc9lhz.gif

OpenStudy (compassionate):

^ Lol. Didn't click.

OpenStudy (anonymous):

Nothing risque or of that sort ;)

OpenStudy (compassionate):

It seems you've convinced me to click your nonsensical link.

OpenStudy (jennychan12):

holy crap look at all these people viewing this...

OpenStudy (help!!!!):

they both suck

OpenStudy (compassionate):

We are conversing, aren't me?

OpenStudy (anonymous):

Yes, yes me am.

OpenStudy (compassionate):

The answer is 6. Deal with it.

OpenStudy (anonymous):

My gifs went to sleep. In other words, too lazy to argue.

OpenStudy (compassionate):

In others words, you lose.

OpenStudy (anonymous):

9

OpenStudy (compassionate):

6

OpenStudy (compassionate):

Wolfram is obviously wrong.

OpenStudy (anonymous):

@Compassionate i'm having trouble understanding what your trying to achieve by being difficult...were you not loved as a child?

OpenStudy (n00bstyle):

So...to conclude: if a person asks this or writes this down, he has to formulate it better! Use another set of brackets.

OpenStudy (compassionate):

I was molested by a reindeer. Don't bring up the bad memories. It's 6.

OpenStudy (anonymous):

^ http://media.tumblr.com/tumblr_mf3rydieE61rbmcbo.gif

OpenStudy (compassionate):

^ http://static.fjcdn.com/pictures/Got_47e91d_1571513.jpg

OpenStudy (kainui):

I kinda didn't like compassionate, but now I actually don't mind him cause he's actually fairly mediocre at trollin'.

OpenStudy (compassionate):

Prove that it's 9. http://i3.kym-cdn.com/photos/images/original/000/166/212/DO%20IT%20owletGOT%202.jpg

OpenStudy (anonymous):

Yeah, trolling isn't really trolling when it's transparent

OpenStudy (kainui):

Unless I'm also trollin'.

OpenStudy (compassionate):

ITT: Trolls trolling trolls who were originally meant to be trolled by trolls.

OpenStudy (kainui):

d'ah!

OpenStudy (anonymous):

Zzzzz

OpenStudy (compassionate):

Cary on my wayward son.

OpenStudy (anonymous):

http://media.tumblr.com/aead2aa43a748ed4d1ed60f9eb7549a6/tumblr_inline_mfjeupVtMI1r9naon.gif I hope there are no children here.

OpenStudy (jennychan12):

the "sqiggles" are integrals :) i don't even know what he did... some math magic wala!

OpenStudy (compassionate):

Rage.

OpenStudy (jennychan12):

@karatechopper the "sqiggles" are integrals :) i don't even know what he did... some math magic wala!

OpenStudy (anonymous):

Ohyes

OpenStudy (karatechopper):

I know they are integrals. but i like to call them squiggles

OpenStudy (jennychan12):

haha oh sowwy but i really don't know what he did. i would say.... BS!!!!

OpenStudy (compassionate):

We can call them squishies.

OpenStudy (abb0t):

squiggles!!!

OpenStudy (kainui):

\[\int\limits_{\int\limits_{\int\limits_{\int\limits_{}^{}}^{\int\limits_{}^{}}}^{\int\limits_{\int\limits_{}^{}}^{\int\limits_{}^{}}}}^{\int\limits_{\int\limits_{\int\limits_{}^{}}^{\int\limits_{}^{}}}^{\int\limits_{\int\limits_{}^{}}^{\int\limits_{}^{}}}} \]

OpenStudy (anonymous):

all this is is a poorly defined statement, but the convention is to go from left to right with MD

OpenStudy (anonymous):

Refer to the following with regard to operator precedence: http://en.cppreference.com/w/cpp/language/operator_precedence 6÷2(1+2) -> (6÷2)*(1+2) -> (3)*(1+2) ->3+6 -> 9